OpenStudy (anonymous):

If the circle (x + 1)2 + (y - 3)2 = 56 is translated 1 units down and 1 units right, then the sum of the new y-intercepts is

OpenStudy (anonymous):

@saifoo.khan can u help me

OpenStudy (anonymous):

btw the 2 after the brackets are menats to be squared. like (x+1) to the power of 2

OpenStudy (anonymous):

the original eq gives a center at (-1,3) . the new translated center is at (0,2) the new eq is x^2+(y-2)^2=56 the y intercepts occur at x=0 so (y-2)^2=56 so y^2-4y+2=56 so y^2-4y-54=0

OpenStudy (anonymous):

using the quadratic formula, we get 4+or- sqrt 232/2 or 4+15.231/2 and 4-15.231/2

OpenStudy (anonymous):

if you add add these up, you get 4

OpenStudy (anonymous):

thnks alot. do u mind if i ask two more questions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

was the answer right?

OpenStudy (anonymous):

Determine the equation of the circle with endpoints of a diameter at (1, 4) and (-3, -6).

OpenStudy (anonymous):

midpoint formula-ceter is at the midpoint os a diameter (1+(-3)/2 and 4-6/2-------->center is (-1,-1) so eq is (x+1)^2+(y+1)^2=r^2 r-- use distance formula ------ sqrt of (2^2+5^2)= sqrt of 29 ----so r^2 =29 got it